OpenStudy (anonymous):

*********************MEDAL AND FAN******************** Find parametric equations for the line which passes through P = (-8, 9, 4) and Q = (8, -6, -5)

OpenStudy (anonymous):

It is a vector of 69

OpenStudy (zzr0ck3r):

@Shrekinstine that makes no sense...

OpenStudy (zzr0ck3r):

we have \(p+t(q-p)\)

OpenStudy (zzr0ck3r):

so you have \(<-8,9,4>+t(<8,-6,-5>-<-8,9,4>) = <-8,9,4>+t<16,-15,-9>\)

OpenStudy (zzr0ck3r):

with me?

OpenStudy (zzr0ck3r):

\(<-8,9,4>+t(<8,-6,-5>-<-8,9,4>) =\\ <-8,9,4>+t<16,-15,-9>\)

OpenStudy (anonymous):

F= a cost +a sin t + bt x = a cos(t), y = a sin(t), z = bt, and you have 2 equation with 2 unknowns. http://en.wikipedia.org/wiki/Parametric_equation#3D_examples

OpenStudy (zzr0ck3r):

this is 3d

OpenStudy (zzr0ck3r):

@TomShoe do you follow?

OpenStudy (zzr0ck3r):

\(<x,y,z>=<-8,9,4>+t<16,-15,-9>\) so we have \(x=-8+16t\\y=9-15t\\z=4-9t\)

OpenStudy (anonymous):

@zzr0ck3r Mostly. You found a formula and are just plugging the values in. That's good so far, now I need to find x= , y=, and z=.

OpenStudy (zzr0ck3r):

\(\uparrow\)

OpenStudy (anonymous):

Did you have to find the magnitude?

OpenStudy (anonymous):

@zzr0ck3r How did you convert each point into an equation with the variable t?

OpenStudy (zzr0ck3r):

\(<\color{blue}{x},\color{red}{y},\color{green}{z}>=<\color{blue}{-8},\color{red}{9},\color{green}{4}>+t<\color{blue}{16},\color{red}{-15},\color{green}{-9}>\)

OpenStudy (zzr0ck3r):

do you see?

OpenStudy (zzr0ck3r):

\(\color{blue}{x = -8+16t}\)

OpenStudy (anonymous):

Oh, I see. Thanks! Sorry for bugging you, just making sure I won't have to ask again :P

OpenStudy (zzr0ck3r):

i much prefer people like you:)